b. The rate law for the reaction (2A+B+C) is -ra=KACA CB, with ka=25 (dm /mol) /s. What are kB and kc?
Answer #1

for the above elementary reaction, the rate of reactions with respect to A, B and C are:

-r_A=K_A\ C_A^2\ C_B\\\\ -r_B=K_B\ C_A^2\ C_B\\\\ r_C=K_C\ C_A^2\ C_B\\\\

and from the elementary reaction given, the rate of reactions are related as:

\frac{-r_A}{2}=-r_B=r_C

hence we get:

\frac{K_A}{2}=K_B=K_C

substituting the values:

K_B=\frac{K_A}{2}\\\\ \Rightarrow K_B=\frac{25\ (dm^3/mol)^2/s}{2}\\\\ \Rightarrow K_B=12.5\ (dm^3/mol)^2/s

and,

K_C=K_B=12.5\ (dm^3/mol)^2/s

hence,

K_B=12.5\ (dm^3/mol)^2/s\\\\ K_C=12.5\ (dm^3/mol)^2/s\\\\